# Nuclear masses and fission barriers within the isospin-square liquid drop model

### v1

@@ IV Potential energy surfaces and fission barrier heights
Figure 2: Macroscopic energy surfaces of 250Cf on the ($c,a_{4}$) plane (top) and around the scission configuration (bottom) are evaluated using the LSD (l.h.s) c and ISLD(r.h.s.) formulae. Here $c$ is the elongation of the nucleus, $a_{4}$ is related to the neck size, and ${\rm A_{h}}$ is the heavy fission fragment mass number PN23 .
The LSD and ISLD macroscopic energy surfaces of 250Cf are compared in Fig. 2. We have used here the FoS shape parametrization PN23 . One has assumed here that the macroscopic energy of a spherical nucleus ($c=1,\,a_{4}=0$) is zero. It is seen that the surfaces obtained in both models are close to each other, and only tiny differences can be noticed. Namely, the ISLD fission valley (top row) is slightly deeper and corresponds to more elongated shapes than the LSD one. The bottom of the LSD valley around the scission configuration (bottom raw) is located at the elongation $c=2.62$. It is located around 4 MeV above the ISLD exit from the valley which appears at $c=2.70$. Such differences in energy and elongation of the nucleus in the macroscopic scission point may have some influence on the total kinetic energy (TKE) of the fission fragments. On the other hand, the stiffnesses of LSD and ISLD potentials with respect to the fission fragment mass asymmetry ($A_{h}$) (see the maps in the bottom raw) are very close to each other. The saddle points in both ($c,\,a_{4}$) maps correspond to almost the same energy ($E_{\rm sadd}\approx 1.3$ MeV and are located around the deformation $c\approx 1.35$ and $a_{4}\approx 0.06$.
@@
Figure 3: Fission barrier heights evaluated with the ISLD set of parameters are compared with the empirical/experimental data taken from Ref. Sm93 (triangles) and Ca09 (stars) and with the theoretical values (squares) obtained within the mac-mic model in Ref. JK21 .
The fission barrier heights estimated using Eq. (5) and the ISLD and the LSD parameters are compared in Fig. 3 with the experimental data taken from Refs. Sm93 ; Ca09 as well as with the theoretical estimates of the highest fission barrier made in Ref. JK21 . The experimental, or better to say empirical, fission barrier heights come mainly from analysis of the fission cross-section energy dependence and fissionability of nuclei, supplemented by data obtained from analysis of the excitation functions of spontaneously fissioning isomers and the group of strong resonances in the sub-barrier fission cross-section Sm93 ; Ca09 . The theoretical barrier heights tabulated in Ref. JK21 were evaluated within the 7D mac-mic model with the YpE macroscopic energy part and the microscopic energy evaluated using the Woods-Saxon (WS) single-particle potential.
As one can see, the fission barrier heights obtained using the Myers and Świa̧tecki topographical theorem and the LSD and ISLD models underestimate, in most cases, the experimental barrier heights while those of Ref. JK21 are, as a rule, larger than the experimental values. One can expect that the fission barrier evaluated in similar as in Ref. JK21 but with the ISLD macroscopic part of the energy will be closer to the data as the topographical theorem says that the effect of the microscopic energy part at the fission saddle is small but not equal to zero.
@@ V Conclusions
We have shown that the new liquid drop mass formula with the asymmetry term proportional to the isospin square (ISLD) describes the presently known experimental and estimates masses well. One has to stress here that the ISLD formula contains only six adjustable parameters. The other models contain more free parameters, e.g., the 21 years old LSD mass formula PD03 , which reproduces the binding energies even better than the FRDM theory MS16 , possesses eight directly fitted parameters and five others (in the congruence/Wigner and odd-even terms) taken from the adjustments made in Ref. MN95 . In this place, it is good to remember that microscopic models, which typically have up to 14 parameters in the interactions, can reproduce 2149 experimental masses with the r.m.s. deviation equal to 0.798 MeV GH09 .
It was also shown that both ISLD and LSD models describe well the fission barrier heights of the heavy nuclei. The ISLD model predicts slightly larger atomic masses than the LSD one in the super-heavy region of nuclei and for neutron-rich isotopes close to the neutron drip line. This difference in the mass estimates could be significant for SHN physics, and it can influence the nuclear r-process probabilities, which is very important in astrophysical theories. Also, the ISLD model predicts more elongated shape of fissioning nuclei at scission configuration than the LSD formula. This effect can influence on estimates of the fission fragment TKE and their charge equilibration (confer Ref. PN24 .
We plan to perform extended dynamical calculations like those made in Refs. PN23 ; PN24 ; PB21 ; KD21 ; LC21 but use the new ISLD formula when evaluating the potential energy surfaces of fissioning nuclei.

### v2

@@ IV Potential energy surfaces and fission barrier heights
The LSD and ISLD macroscopic energy surfaces of 250Cf are compared in Fig. 3. We have used here the FoS shape parametrization PN23 . It was also assumed that the macroscopic energy of a spherical nucleus ($c=1,\,a_{4}=0$) is zero. It is seen that the surfaces obtained in both models are close to each other, and only tiny differences can be noticed. Namely, the ISLD fission valley (top row) is slightly deeper and corresponds to more elongated shapes than the LSD one. The exit of the LSD valley around the scission configuration (bottom raw) is located at the elongation $c=2.62$. It is lying around 4 MeV above the ISLD exit, which appears at $c=2.70$. Such differences in energy and elongation of the nucleus in the macroscopic scission point may have some influence on the total kinetic energy (TKE) of the fission fragments. On the other hand, the stiffnesses of LSD and ISLD potentials with respect to the fission fragment mass asymmetry ($A_{h}$) (see the maps in the bottom raw) are very close to each other. The saddle points in both ($c,\,a_{4}$) maps correspond to almost the same energy ($E_{\rm sadd}\approx 1.3$ MeV and are located around the deformation $c\approx 1.35$ and $a_{4}\approx 0.06$.
Figure 4: Fission barrier heights evaluated using the topographical theorem MS96 with the ISLD(a) (crosses), ISLD(b) (points), and LSD (open circles) set of parameters are compared with the empirical/experimental data taken from Ref. Sm93 (triangles) and Ca09 (stars) and with the theoretical values (squares) obtained within the mac-mic model in Ref. JK21 .
@@
The fission barrier heights estimated using Eq. (6) and the ISLD (a) (crosses) and (b) (circles) and the LSD (open circles) formulas are compared in Fig. 4 with the experimental data taken from Refs. Sm93 (triangles) and Ca09 (stars). In addition, the theoretical estimates of the highest fission barrier done in Ref. JK21 are marked by squares. The experimental, or better to say empirical, fission barrier heights come mainly from analysis of the fission cross-section energy dependence and fissionability of nuclei. They are supplemented by data obtained from analysis of the excitation functions of spontaneously fissioning isomers and the group of strong resonances in the sub-barrier fission cross-section Sm93 ; Ca09 . The theoretical barrier heights tabulated in Ref. JK21 were evaluated within the 7D mac-mic model with the YpE macroscopic energy part and the microscopic energy evaluated using the Woods-Saxon (WS) single-particle potential.
As one can see, the fission barrier heights obtained using the Myers and Świa̧tecki topographical theorem and the LSD and ISLD models underestimate, in most cases, the experimental barrier heights while those of Ref. JK21 are, as a rule, larger than the experimental values. The ISLD barriers obtained with the set (b) parameters are very close to the LSD ones, while those computed with the set (a) are, on average, by 1 MeV smaller. One can expect that the fission barriers evaluated in a similar as in Ref. JK21 way, but with the ISLD macroscopic part of the energy will be closer to the data as the topographical theorem estimates the barrier heights from below MS96 .
@@ V Conclusions
We have shown that the new liquid drop mass formula with the charge asymmetry term proportional to the isospin square (ISLD) describes the presently known experimental and estimated from systematic atomic masses well. One has to stress here that the ISLD formula contains only six adjustable parameters. The other models contain more free parameters, e.g., the twenty-one years old LSD mass formula PD03 , which reproduces the binding energies even better than the FRDM theory MS16 , possesses eight directly fitted parameters and four others (in the congruence/Wigner and odd-even terms) taken from the adjustments made in Ref. MN95 . In this place, it is good to remember that microscopic models, typically heaving up to 14 interaction parameters, can reproduce 2149 experimental masses with the r.m.s. deviation equal to 0.798 MeV GH09 .
It was found that all experimentally estimated masses of superheavy nuclei with a neutron number larger than 160 have up to 3 MeV larger masses than the predicted ones. What is the origin of these discrepancies: inaccuracies in evaluating the microscopic energy, wrong asymptotic behavior of our macroscopic models, or inaccurate evaluation of the atomic masses from the systematics? More detailed calculations to answer this question are necessary.
It was also shown that both ISLD (b) and LSD models describe well the fission barrier heights of the heavy nuclei. The ISLD model predicts slightly larger atomic masses than the LSD one in the superheavy region of nuclei and for neutron-rich isotopes close to the neutron drip line. This difference in the mass estimates could be significant for SHN physics, and it can influence the nuclear r-process probabilities, which is very important in astrophysical theories. Also, the ISLD model predicts a more elongated shape of fissioning nuclei at scission configuration than the LSD formula. This effect can influence estimates of the fission fragment TKE and their charge equilibration (confer Ref. PN24 ).
We plan to perform extended dynamical calculations like those made in Refs. PN23 ; PN24 ; PB21 ; KD21 ; LC21 but the use of the new ISLD formula when evaluating the potential energy surfaces of fissioning nuclei. The well-defined deformation dependence of the macroscopic energy term linear in $|N-Z|$ may also influence the isotopic yields of the fission fragments (confer Ref. PN24 ).
Acknowledgements
The author thanks Kai Neergård for his helpful remarks.